OpenStudy (anonymous):

Frank kicks a soccer ball off the ground and in the air with an initial velocity of 30 feet per second. Using the formula H(t) = −16t2 + vt + s, what is the maximum height the soccer ball reaches?

OpenStudy (anonymous):

14.1 feet 13.7 feet 13.2 feet 15.2 feet

OpenStudy (anonymous):

\[v=30,s=0\] so the equation is \[h(t)=-16t^2+30t\]

OpenStudy (anonymous):

max is at the vertex first coordinate of the vertex is \(-\frac{b}{2a}\) which in your case is \[-\frac{30}{2\times -16}=\frac{30}{32}=\frac{15}{16}\]

OpenStudy (anonymous):

or if you perfere \(.9375\) for the second coordinate of the vertex, which is what you are looking for, compute \[-16(.9375)^2+30\times .9375\] with a calculator

OpenStudy (anonymous):

so what is it

OpenStudy (anonymous):

14.1??:)

OpenStudy (anonymous):

yeah that is the closest